# Photoelectron Circular Dichroism of Aqueous-Phase Alanine

### v1

@@ 2 Results and discussion
Following the process outlined in the supporting information and using Eq. 2, we determined a single value of $b^{+1}_{1}$ per carbon center for each pair of $+$ and -- measurements for a given enantiomer, solution pH, and photon energy (Fig. 3). The data were then averaged within 250 meV KE windows for visualization. The complete datasets without averaging are presented in the supporting information (Figs. LABEL:SI-fig:S4, LABEL:SI-fig:S8, LABEL:SI-fig:S12). The analysis was carried out blind with respect to the enantiomer, such that we analyzed all the data without knowing whether any given pair of measurements corresponded to a solution of L-, D-, or DL-alanine. An overview of the $b^{+1}_{1}$ parameters obtained for C1 at different pH values is presented in Fig. 4. Beginning with the pH 1 data, corresponding to $Ala^{+}$, weak enantiomer-dependent asymmetry in $b^{+1}_{1}$ may be seen for photoelectrons with KE 9 eV. At this energy, L-alanine exhibit a negative value of $b^{+1}_{1}$, while D-alanine yields a somewhat positive value. These values appear to rapidly trend toward zero at higher kinetic energies, and no significant PECD effect is found above 10 eV. For $Ala^{zw}$ at pH 7, the situation is different. Here, $b^{+1}_{1}$ is close to zero in the KE range of 9–11 eV, but appears to increase in magnitude briefly around 12–13 eV. From 14–16 eV, the data is noisy and does not indicate clear PECD. At 17 eV, $b^{+1}_{1}$ increases once more. The sign of $b^{+1}_{1}$ is the same for the corresponding enantiomer at pH 1. The clearest PECD can be seen for the case of $Ala^{-}$. For this protonation state, $b^{+1}_{1}$ as measured for C1 is clearly non-zero for the entire KE range measured, even up to 17 eV. The sign of $b^{+1}_{1}$ remains constant throughout the entire range, and is consistent with that observed for $Ala^{zw}$ and $Ala^{+}$.
@@
Figure 4: Values of the $b_{1}^{+1}$ photoionization parameter obtained for C 1s measurements of aqueous solutions of D-, L-, and DL-alanine (red, blue, and green points, respectively) at pH 1, 7, and 13 (top, middle, and bottom; corresponding to the cationic, zwitterionic, and anionic form of the molecule, respectively). All $b_{1}^{+1}$ values shown correspond to photoionization of the C1 carboxylic group. The data displayed is the result of binning the results of individual measurement sets using a kinetic-energy window of 250 meV. The shaded error bars represent the combined error of the binned data points. A description of the error propagation may be found in the supporting information.
@@
We were unable to conclusively identify signatures of PECD for the C2 and C3 atoms, with the possible exception of $Ala^{zw}$ in the KE range of 15–19 eV (Figs. LABEL:SI-fig:S7 and LABEL:SI-fig:S8). However, due to the limited quantity of data in this KE range for $Ala^{zw}$ and the overlapping nature of these peaks, we do not consider the C2 and C3 data to convincingly demonstrate PECD. As mentioned above, we dedicated the most time to collecting data points for $Ala^{-}$, which appeared from the onset to yield the most clear PECD. As such, we are the most confident in our results for $Ala^{-}.$ Although it may seem surprising that C1, C2, and C3 should exhibit such large differences in PECD magnitude, significant site-dependent PECD has previously been reported in C 1s studies of carvone and fenchone,49, 14 thereby highlighting the sensitivity of PECD to the atomic site probed. The datasets corresponding to Fig.4 for C2 and C3 are available in the supporting information (see Figs. LABEL:SI-fig:S7 and LABEL:SI-fig:S11).
Taken altogether, the data reveal a significant PECD effect for aqueous-phase alanine in photoionization of the C1 carbon group, for all three of alanine’s protonation states. The sign of $b^{+1}_{1}$ is consistent between $Ala^{+}$, $Ala^{zw}$, and $Ala^{-}$, but differences in magnitude are apparent, with the most prominent effects measured for $Ala^{-}$. In contrast to most other PECD studies, we do not note a clear KE dependence for $b^{+1}_{1}$ in our data. This is likely due to the compensation of increasing PECD by concomitantly increasing elastic scattering at low KEs (as illustrated in Fig. 2 within the blue and green regions). The extent to which the measured $b^{+1}_{1}$ values will be reduced by scattering will depend on the specific KE-dependent elastic-scattering cross section. As such, substantial electron scattering may result in changes not only to the magnitude of the measured $b^{+1}_{1}$ values, but also their apparent KE dependency. This result suggests that, for the current case of alanine in water, there may be limited experimental benefit (in terms of magnitude of measured $b^{+1}_{1}$) in focusing on very-low KE photoelectrons. As these low-KE spectra are the most difficult to analyze due to the necessity for significant background subtraction, follow-up studies may benefit from investigation of higher-energy (KE = 15–25 eV) photoelectrons. We can estimate the average degree of $b^{+1}_{1}$ attenuation in our present measurements through comparison with previous work. As mentioned above, the angular asymmetry of the O 1s PAD is reduced by 60% for liquid compared to gaseous water at KE = 25 eV 34. Similarly, in our previous core-level PECD study of liquid fenchone 30, we reported a factor of 5 decrease for $b^{+1}_{1}$ compared to that measured for gas-phase fenchone within the KE range of 9–13 eV 14. This seems to be a general trend upon condensation of chiral molecules; the magnitude of PECD measured for free serine in the gas phase was also found to decrease upon condensation of serine into nanoparticles, although in this case condensation was also accompanied by changes of molecular charge state and conformer population, both of which influence PECD in a non-trivial manner 18. A more detailed understanding of elastic scattering of low-energy photoelectrons in aqueous solutions of alanine will be required to more accurately characterize the degree of isotropization in the present study. However, based on the work mentioned above, we estimate that our present results for $b^{+1}_{1}$ may be attenuated by a factor of 3–5. For the C1 data shown in Fig. 4, this means that our measured values of $b^{+1}_{1}$approx 0.02 likely correspond to initial $b^{+1}_{1}$ values of approximately 0.06–0.1 prior to deleterious photoelectron scattering. The above arguments equally hold for the C2 and C3 carbons and would serve to make what may have been initially low values $b^{+1}_{1}$ for those atoms undetectable with the current experimental sensitivity.
Although no other core-level PECD studies of alanine have been published, we may compare our results with the existing theoretical and experimental investigations of valence-band PECD of alanine. These studies have revealed only modest PECD, with $b^{+1}_{1}~{}\approx~{}0.02$, corresponding to a PECD effect of 4% 17, 50. This value is roughly 10 times larger than the comparable ECD magnitude (g factor) for alanine in the gas phase 51, but significantly lower than the asymmetry reported for valence-band photoemission from rigid molecules such as terpenes ($b^{+1}_{1}$ up to $\approx$ 0.1)52. The relatively weak PECD reported for alanine is similar to that observed for other flexible molecules in the gas phase, including alaninol 53, indanol 54, 37, 22, endoborneol 55, and proline 19, and may have to do with conformational averaging. Reflecting this conformational sensitivity, most PECD studies in the literature focus on relatively rigid molecules, including bicyclic terpenes and terpenoids 4, 11, 56, 12, 13, 15, 52, 57, 30, or other heterocylic oxirane derivatives 58, 59, 60, 8, 61, 62, 10.
If core-level photoionization of alanine’s C2 and C3 atoms in the gas phase were to result in similar values as measured for valence-band photoionization of alanine ($b^{+1}_{1}$approx 0.02). they would likely manifest in our experiments as $b^{+1}_{1}$approx 0.004-0.007 (based on our scattering estimate above). As can be seen in Fig. 4, the conclusive detection of such low values is currently not feasible. Considering the general trend of flexible molecules exhibiting weaker PECD, it appears reasonable that we currently lack the sensitivity to PECD from all of alanine’s carbon groups. Somewhat more surprising is that, despite elastic scattering, our measured values of $b^{1}_{1}$ for core-level photoionization of C1 in aqueous $Ala^{-}$ are comparable to those reported for valence-band photoionization of gas-phase $Ala^{0}$.
The specific details, signs, and magnitudes of PECD arise as results of extensive interference between multiple continuum partial electron waves upon photoionization, and thus defy intuitive prediction 47. Nevertheless, it is clear that the local scattering environment must necessarily influence the continuum wave function of each partial electron wave. The significant role of final-state scattering in core-level PECD was first experimentally demonstrated in C 1s photoionization of gaseous carvone and fenchone, for which strong PECD was observed despite the spherical symmetry of the initial states 49, 14. As our present study constitutes the only report of PECD of aqueous-phase systems at this time, it is useful to briefly discuss how some of the most prominent differences between aqueous-phase and gas-phase molecules may manifest in PECD measurements. Theoretical approaches to determining PECD are currently limited to small gas-phase molecules, and we hope that the following discussion will stimulate the development of new methods capable of modeling PECD for aqueous-phase systems, which we anticipate will prove invaluable for the the application of liquid-jet PECD as a practical analytical technique. We suggest that a model of PECD for zwitterionic alanine microsolvated by 5 explicit water molecules — the minimum required to stabilize the charged forms of alanine in solution 63 — would constitute a valuable step in this direction.
@@
The most energetically favorable conformations of alanine in the gas phase are stabilized by intramolecular hydrogen-bond formation between the NH2 and COOH groups 65, 66, 67. These intramolecular hydrogen bonds will likely be disrupted in solution, being replaced by intermolecular alanine-solvent interactions. The loss of intramolecular hydrogen bonds may result in subtle changes to PECD, as intramolecular hydrogen-bond orbital polarization has been shown to play a small yet significant role in determining chiral PAD asymmetry 68. The relevance of intermolecular interactions for PECD has been revealed through studies of homochiral clusters of small, weakly interacting molecules, for which profound changes in PECD were observed between isolated monomers, dimers, and larger "n-mers" 69, 16. In the present context, the formation of new intermolecular hydrogen bonds between water and alanine will almost certainly substantially influence the measured PECD. We also note that a molecule’s lowest-energy conformers well may differ between the gas and solution phases 46, and that such differences would be associated with changes in the nuclear scattering potential and the consequent PECD signal.
Although only little work focused on the solvation environment of alanine in solution across its different protonation states exists, we will attempt to sketch a preliminary model based on the current state of the literature. In the case of $Ala^{zw}$, water molecules readily hydrogen bond to the charged COO- and NH+${}_{3}^{+}$ groups, forming a bridge between them and thus stabilizing the doubly ionized form of the molecule 70, 71, 63, 72. At this pH, water molecules slightly prefer bonding to COO-, forming on average six bonds with that group compared to three with NH+${}_{3}^{+}$ 73. This preference of water to form hydrogen bonds with the COO- group is consistent for $Ala^{-}$ and is also similar to what has been previously reported for zwitterionic glycine in water 74. Under acidic conditions, the trend seems to invert, with four waters preferentially coordinating to the NH+${}_{3}^{+}$ group and only two to the COOH group 73. Simulations of alanine in the presence of larger numbers of explicit water molecules have demonstrated that alanine–water complexes are dynamic, with alanine’s different functional groups capable of rotating, and with hydrogen-bonded water molecules within the first solvation shell constantly exchanging with those from the second 70. However, these processes occur on the ps scale, and while they may be significant for time-resolved PECD studies, alanine’s solvation shell appears relatively structured and rigid averaged over the time scale of minutes, with the charged functional groups hydrogen-bonded to water molecules more than 95% of the time 70, 71. Even alanine’s methyl group, which does not form strong hydrogen bonds with water, seems to be involved in the formation of a structured screening water network 75 These studies identify large differences in the degree of hydrogen bonding between water molecules and alanine’s different functional groups depending on pH, which may partially explain the site- and protonation-state-dependent PECD that we have measured.
PECD has recently been reported for gas-phase photoionization of an achiral chromophore weakly bound to a chiral molecule, indicating that PECD exhibits long-range (on the 5 Å scale) sensitivity to chirality 76. As such, PECD of aqueous-phase chiral molecules is almost certain to be sensitive to induced solvent chirality within the first solvation shell. The extent to which induced chirality may arise in a solvation shell appears to depend strongly on solvent polarizability and the van der Waals volume of the solvent molecule 28, 77, 78. Chiral sum frequency generation (SFG) spectroscopy experiments probing induced chirality in the solvation shell around small proteins near the liquid-vapor interface have revealed that the most significant chiral signal from the solvent originates within the first solvation shell, with the second and third solvation shells being nearly achiral 79. These conclusions are in agreement with results from molecular dynamics simulations 78 and suggest an analogy of molecules in the first solvation behaving cooperatively as an achiral glove (for example, a non-handed nitrile glove), adopting a preferential handedness only upon being worn (or upon solvating a chiral solute). The consequences of induced chirality for chiral spectroscopy can be significant, as in the case of methyloxirane in benzene, for which the ECD signal appears to be dominated by the solvation-shell structure rather than by the intrinsically chiral solute itself 80. Vibrational circular dichroism (VCD), probing the difference in absorption of left- and right-handed CPL in the infrared and near-infrared region, has also proven capable of resolving signatures of chirality transfer between solute and solvent via hydrogen bonding 81, 82, 83, 84. Although the chirality of alanine’s aqueous solvation shell has not yet been studied directly, all criteria seem to be met for the formation of a chiral solvation shell. We note that the ECD response of $Ala^{zw}$ differs substantially in the solid vs. in the liquid phase 85 and in general the ECD spectra of aqueous-phase alanine is dependent on solution pH 3, but it is presently unclear whether these trends are related to solvent effects, or simply due to differences in solute conformation or protonation state. We suggest that molecular dynamics studies investigating the chirality of aqueous-phase alanine’s solvation shell as a function of protonation state, combined with new approaches to calculating PECD for larger systems, will be of direct relevance for interpreting our present experimental results. The chirality of a solvation shell will likely manifest differently for different chiral spectroscopic techniques. ECD and VCD appear capable of revealing general chiral effects in the solvation shell, and chiral SFG is well suited for studying molecules at the liquid-vapor interface. Due to its site specificity, core-level PECD may enable much more localized probing of chiral solvation effects, both for bulk-soluble and surface-active molecules. This study constitutes an important first step in this direction.
Importantly, we note that although PECD of aqueous-phase chiral molecules currently requires time-intensive measurement, new experimental approaches that enable simultaneous detection of photoelectrons in both the forward and backward directions with enhanced detection efficiency would dramatically expand the general applicability of this method. In particular, we highlight liquid-compatible velocity map imaging as a technique of considerable interest for the further development of LJ-PECD 86.
3 Conclusions
We have measured core-level photoelectron circular dichroism (PECD) for aqueous-phase alanine, the smallest chiral amino acid, for each of its three charge states (cationic, zwitterionic, and anionic). We detected significant PECD for photoionization of alanine’s carboxylic acid functional group, which exhibited a clear magnitude dependence across pH conditions, but did not exhibit a strong kinetic-energy dependence. We attribute the latter observation to significant kinetic-energy dependent elastic scattering of low-energy primary photoelectrons. We did not measure significant PECD in photoionization of alanine’s chiral center or methyl group, suggesting that the effects for those groups are weak and below our currently modest experimental sensitivity. For photoionization of the carboxylic acid, PECD is the most pronounced under basic solution conditions, where alanine adopts the anionic form, decreases for the zwitterionic form, and is very weak for the cationic form. The differences observed across protonation states may be due to a number of factors, including changes in molecular charge state, functional-group dependent solute-solvent interactions, and first solvation-shell structure. Advances in computational approaches for modeling PECD in multiple-molecule systems will be necessary to identify the relative weights of these effects, and higher collection efficiency detectors capable of simultaneous measurement in the forward and backward direction will dramatically simplify its application in the future. Our demonstration of aqueous-phase PECD marks a significant advance in the field of liquid-jet photoelectron spectroscopy. Given its chemical-state and site specificity and enantiomeric sensitivity, liquid-phase PECD is likely to become a remarkably useful analytical method for studying chiral molecules under biologically relevant aqueous conditions.
4 Experimental
We carried out all aqueous-phase PECD experiments using our custom LJ-PES apparatus — Electronic Structure from Aqueous Solutions and Interfaces (EASI) — developed specifically for this purpose 29. We introduced alanine aqueous solutions into the experimental chamber using a custom quartz capillary, with an inner diameter on the order of 20–30 $\mu$m. Solutions were driven by a high-performance liquid chromatography pump (Shimadzu LC–20AD) at a flow rate of 0.8–1.5 mL/minute, resulting in solution velocities of 20–80 m/s. The solutions were prepared by diluting alanine in water, adjusting solution pH where necessary with concentrated HCl or solid NaOH, and finally diluting the solution to reach a final concentration of 1 M alanine. The solutions were injected horizontally and frozen out after the interaction region via contact with a liquid-nitrogen-cooled cold trap. All PES experiments were performed at the P04 soft x-ray beamline of the PETRA III storage ring at DESY in Hamburg, Germany. Critically, the P04 beamline is equipped with an APPLE-II undulator, enabling experiments with circularly polarized light within the experimentally relevant photon-energy range of 290-320 eV 87. Although not available during our early measurements, recent advancements at the P04 beamline now enable the estimation of the degree of circular polarization for a given set of undulator conditions. Retroactive estimates revealed an expected circular polarization of >99% for all experimental conditions employed in this study. D-, L-, and DL-alanine (>98% purity) were purchased from Sigma-Aldrich and Carl Roth and used as received. Enantiomeric purity was confirmed via intermittent testing using ECD (see Figure LABEL:SI-fig:S1). For additional experimental details regarding our instrumentation or LJ-PES more generally, we direct interested readers to our recent technical publication 29. All data reduction was performed using the Igor Pro analysis software (Wavemetrics,version 9).
@@ 4 Experimental
We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association (HGF), for the provision of experimental facilities. Parts of this research were carried out at PETRA III at the P04 soft x-ray beamline. Beamtime was allocated for proposals II-20180012, I-20200682, II-20210015, I-20211126, and I-20230378. We thank the P04 beamline staff, in particular Dr. Moritz Hoesch and Jörn Seltmann, for their long-standing support for this project. We also thank the European Molecular Biology Laboratory for access to the circular dichroism spectrophotometer. We thank Prof. Petr Slavíček, Prof. Philipp Demekhin, former group members Dr. Anne Stephansen, Dr. Claudia Kolbeck, Dr. Marvin Pohl, Dr. Karen Mudryk, and Dr. Tillmann Buttersack, as well as current group members Harmanjot Kaur, Qi Zhou, Dr. Lukáš Tomaník, Dr. Nicolas Velasquez, Henrik Haak, and Sebastian Kray for valuable discussion and assistance with the transportation and preparation of our setup (EASI) for measurements at PETRA III.
@@
The supporting information contains electronic circular dichroism measurements of aqueous solutions of alanine, calculations of the photoionization parameter $b^{0}_{2}$ in the kinetic-energy range of 0–25 eV, a summary of all values of $b^{+1}_{1}$ measured for alanine’s different carbon groups, including values calculated before background subtraction and kinetic-energy averaging was performed, and a description of the means determining the error of the $b^{+1}_{1}$ values, as shown in Figures 3 and 4.

### v2

@@ 2 Results
Following the process outlined in the supplementary information and using Eq. 2, we determined a single value of $b^{+1}_{1}$ per carbon center for each pair of ++ and - measurements for a given enantiomer, solution pH, and photon energy (Fig. 3). The data were then averaged within 250 meV KE windows for visualization, and to provide a realistic accuracy estimate for our measured photoelectron kinetic energies (see above for factors influencing the accuracy of "as-measured" kinetic energies). The complete datasets without averaging are presented in the supplementary information (Figs. LABEL:SI-fig:S4, LABEL:SI-fig:S8, LABEL:SI-fig:S12). The analysis was carried out blind with respect to the enantiomer, such that we analyzed all the data without knowing whether any given pair of measurements corresponded to a solution of L-, D-, or DL-alanine. An overview of the $b^{+1}_{1}$ parameters obtained for C1 at different pH values is presented in Fig. 4. Beginning with the pH 1 data, corresponding to $Ala^{+}$, weak enantiomer-dependent asymmetry in $b^{+1}_{1}$ may be seen for photoelectrons with KE 9 eV. At this energy, L-alanine exhibit a negative value of $b^{+1}_{1}$, while D-alanine yields a somewhat positive value. These values appear to rapidly trend toward zero at higher kinetic energies, and no significant PECD effect is found above 10 eV. For $Ala^{zw}$ at pH 6, the situation is different. Here, $b^{+1}_{1}$ is close to zero in the KE range of 9–11 eV, but appears to increase in magnitude briefly around 12–13 eV. From 14–16 eV, the data is noisy and does not indicate clear PECD. At 17 eV, $b^{+1}_{1}$ increases once more. The sign of $b^{+1}_{1}$ is the same for the corresponding enantiomer at pH 1. The clearest PECD can be seen for the case of $Ala^{-}$. For this protonation state, $b^{+1}_{1}$ as measured for C1 is clearly non-zero for the entire KE range measured, even up to 17 eV. The sign of $b^{+1}_{1}$ remains constant throughout the entire range, and is consistent with that observed for $Ala^{zw}$ and $Ala^{+}$. Due to the very time-consuming nature of the experiments, it was not always possible to perform the same number of repeat measurements for D-, L-, and racemic alanine solutions. As such, it is critical to keep in mind that the points plotted in Fig. 4 do not represent the same amount of measurements. This may be seen clearly for the unbinned data shown in Fig. LABEL:SI-fig:S4.
@@
Figure 4: Values of the $b_{1}^{+1}$ photoionization parameter obtained for C 1s measurements of aqueous solutions of D-, L-, and DL-alanine (red, blue, and green points, respectively) at pH 1, 7, and 13 (top, middle, and bottom; corresponding to the cationic, zwitterionic, and anionic form of the molecule, respectively). All $b_{1}^{+1}$ values shown correspond to photoionization of the C1 carboxylic group. The data displayed is the result of binning the results of individual measurement sets using a kinetic-energy window of 250 meV. The shaded error bars represent the combined error of the binned data points. A description of the error propagation may be found in the supplementary information.
@@
We were unable to conclusively identify signatures of PECD for the C2 and C3 atoms, with the possible exception of $Ala^{zw}$ in the KE range of 15–19 eV (Figs. LABEL:SI-fig:S7 and LABEL:SI-fig:S8). However, due to the limited quantity of data in this KE range for $Ala^{zw}$ and the overlapping nature of these peaks, we do not consider the C2 and C3 data to convincingly demonstrate PECD. As mentioned above, we dedicated the most time to collecting data points for $Ala^{-}$, which appeared from the onset to yield the most clear PECD. As such, we are the most confident in our results for $Ala^{-}.$ It may seem surprising that C1 exhibits clear PECD, while C2 and C3 do not. However, significant site-dependent PECD has previously been reported in C 1s studies of fenchone and glycidol,14, 51, 52 thereby highlighting the sensitivity of PECD to the specific atomic site probed. The datasets corresponding to Fig.4 for C2 and C3 are available in the supplementary information (see Figs. LABEL:SI-fig:S7 and LABEL:SI-fig:S11).
3 Discussion
These results provide the only currently available report of PECD in aqueous-phase systems. They reveal a significant photoelectron circular dichroism (PECD) effect for aqueous-phase alanine in photoionization of the C1 carbon group which falls within the range 0.5% – 40% commonly encountered in gas-phase PECD studies. The sign of $b_{1}^{+1}$ is consistent across all three protonation states but there are differences in magnitude, with the most pronounced effects seen for $Ala^{-}$. Unlike many gas-phase PECD studies, we do not note a significant variation of the chiral asymmetry, $b_{1}^{+1}$, with photon energy.
Any PECD from the C2 and C3 photoemission essentially must fall below our presently achieved detection limits using a cylindrical liquid jet. Based upon previous experience with this arrangement we can estimate a possible reduction of any photoelectron anisotropy due to electron re-scattering within the bulk aqueous phase. In liquid water the angular anisotropy factor, $\beta$, of the O 1s PAD is reduced by 60% compared to that of gaseous water at KE = 25 eV 34. Similarly, in our previous core-level PECD study of liquid fenchone 30, we reported a factor of 5 decrease for $b_{1}^{+1}$ compared to that measured for gas-phase fenchone within the KE range of 9-13 eV 14. Hence, we tentatively estimate that our present results for $b_{1}^{+1}$ may be attenuated by a factor of 3–5. For the C1 data shown in Fig. 4, this means that our measured values of $b_{1}^{+1}\approx 0.02$ may correspond to initial $b_{1}^{+1}$ values of approximately 0.06–0.1 prior to deleterious photoelectron scattering.
There is, a priori, no particular surprise that the C1 PECD should exceed that of the asymmetrically substituted C2 site – the so-called chiral center. The specific details, signs, and magnitudes of PECD are the result of extensive interference between the many continuum partial electron waves during photoionization, making any intuitive prediction or interpretation of signs and magnitudes extremely challenging. It is also well established that PECD asymmetry results from multiple scattering off the spatially extended chiral molecular framework 11, 14 and so is not solely contributed by a designated "chiral center". Indeed, strong PECD can be observed in molecules with no asymmetrically substituted centers but that display other chiral structural forms such as axial- or propeller-ligand symmetries. 53, 20. Calculations for e.g. fenchone 51 clearly show how the variation in $b_{1}^{+1}$ values over an extended kinetic-energy range mean that the relative signs and magnitudes observed from individual emitter sites will vary with observation energy.
PECD of gas-phase alanine has been reported in the valence region 17, 54 but no C 1s measurements have been published so far. Were they available, simple comparisons would nevertheless be impeded by the different forms (ionic, zwitterionic) that alanine adopts in solution, and that we specifically targeted in this work. Then, even if these charge differences could be incorporated into current modeling, there is in general little value in comparing gas-phase measurements with those made in aqueous media without some understanding of further ways in which solvation may influence the observed PECD. Since our work aims to stimulate the development of theoretical modeling for liquid-phase PECD and, conversely, to show the extensive detail that may then be revealed with the benefit of comprehensive interpretive tools, we next set out further considerations of solvation effects.
The scattering of emitted photoelectrons by bulk water has been alluded to above, but more intimate interactions between a molecule and its solvation shell require consideration. The PECD response is known to be very sensitive to molecular conformation, with great variation in magnitude and sign of the chiral asymmetry even with small changes such as the orientation of a terminal -OH rotamer 52, 55. This can be a hindrance or a benefit depending on one’s perspective and purpose. Solvent–molecule interactions may change the molecular conformational landscape. Approximate treatments for solvation effects have been implemented, and are been widely applied, in many electronic-structure calculation packages. For example, the solvation shell of zwitterionic alanine and its consequent conformational rearrangement have been studied using such methods 56, but while these approaches may help identify preferred new conformations in a polar medium, they will lack explicit treatments of individual solvent molecules and their more specific interactions that can provide the level of insight likely to be required for accurate liquid-phase PECD calculations.
A study of microsolvated structures of protonated alanine 57 details how the addition of solvent molecules leads to the disruption of the intermolecular hydrogen bonds that stabilize and determine gas-phase conformers, replacing them with intramolecular molecule-solvent bonding until the first solvation shell completes with four attached water molecules. Such rearrangement, combined with stabilization of charge in the polar medium, can give rise to the phenomenon of conformational locking in solution, reducing the available conformational space as was inferred for the amino acid proline 45. On the other hand, conformational unlocking going from the gas phase to solution has been reported for asparagine 58, 59, and molecular dynamics simulations of zwitterionic alanine in small water droplets have indicated that alanine does not adopt a single preferred conformer in aqueous solution 60. Either an increase or a decrease in conformation flexibility could result in significant PECD changes, and would hence need careful identification. Even if a molecule were conformationally rigid, and so the chiral nuclear scattering potential relatively unaffected by solvation, there is evidence that PECD may be sensitive to the polarization of the electron density by H-bonding 16, and could thereby change when H-bonding was disrupted by solvation.
Our results indicate that, even for molecules without significant surface activity, surface-orientation effects may not be trivial. Although it is currently not possible to correlate orientation of molecules at the liquid-vapor interfaces to observed PECD, given that orientation appears to increase the magnitude of PECD in gas-phase experiments 8, 9, 10, explicit considerations of both surface propensity and molecular orientation seem to be necessary prerequisites for attempting to model solution-phase PECD.
The central importance of multiple scattering off the chiral molecular potential is particularly evident in core-level PECD when photemission occurs from an essentially spherical – hence achiral – 1s orbital. It is also commonly evident that such chiral final-state scattering may be a long range effect 61. A number of investigations have examined PECD from weakly bound homochiral gas-phase clusters, of species as large as camphor dimers 15, that show changes in chiral angular distribution compared to the monomer. In a study of H-bonded n-mers of glycidol, a total sign inversion of the strong monomer PECD was observed in the clustered species, with less pronounced differences between the individual $n=2,3,4$ clusters 16. More recently, a study of an achiral–chiral complex of phenol-methyloxirane 62 has shown evidence, namely photoelectron chiral asymmetry from ionization of the phenol chromophore, of induced chirality through such loose bonding. Such chirality transfer between solute and achiral solvent is well documented in, for example, vibrational circular dichroism (VCD)63, 64, 65, 66. Hence, it seems inevitable that the detailed structure and interactions of at least the first solvation shell 67, 68 would need to be incorporated into a realistic model of aqueous-phase PECD. In general, it appears likely that any effective model of PECD for solvated species must begin with a highly accurate representation of molecular conformation, hydrogen-bonding configuration, and molecular surface propensity and orientation.
One restriction on the currently achieved experimental sensitivity in this investigation stems from the lengthy data acquisition times per measurement and the consequent limitations on the statistical data quality. This is largely a consequence of the detection arrangement, whereby the electron analysis is performed by a hemispherical analyzer that can only measure at a small fixed solid angle and has to be scanned over the electron energy range of interest. Gas-phase PECD measurements were greatly advanced by the early adoption 12 of velocity map imaging (VMI), which permits energy-resolved photoelectron angular distributions to be detected in a single measurement69. The development of liquid-jet-compatible velocity map imaging is now an ongoing project, which we anticipate will dramatically decrease the amount of time needed for liquid-jet PECD measurements in the future, and thus significantly improve the general suitability of the technique for probing solution-phase molecular chirality.
4 Conclusions
We have measured core-level photoelectron circular dichroism (PECD) for aqueous-phase alanine, the smallest chiral amino acid, and examined each of its three charge states (cationic, zwitterionic, and anionic) by the variation of solution pH. A significant PECD was detected for photoionization of alanine’s carboxylic acid functional group. This exhibited a clear magnitude dependence across pH conditions, but did not exhibit a strong kinetic-energy dependence. On the other hand, neither the photoionization of alanine’s asymmetrically substituted central carbon, nor its methyl carbon, displayed significant PECD, indicating that the effects for those groups are weaker and below our currently modest experimental sensitivity.
For photoionization of the carboxylic C 1s, PECD is the most pronounced under basic solution conditions, where alanine adopts the anionic form, but is very weak for both the zwitterionic and cationic forms. The differences observed across protonation states may be due to a number of factors, including changes in molecular charge state, conformation, surface propensity, molecular orientation, functional-group dependent solute-solvent interactions, and first solvation-shell structure.
Theoretical approaches to PECD are currently limited to small gas-phase systems and we identify some of the factors that may be required to develop realistically complete theoretical models for PECD effects in solution. Conversely, this discussion allows one to identify the scope of additional insight into the solvated state that could be exposed by the interpretation of PECD measurements. Experimentally, feasible technique developments of VMI to provide energy- and angle-multiplexed detection, coupled with the introduction of ultrathin liquid-jet technologies70 makes clear the potential for significantly improved data collection rates and enhanced detection sensitivity.
Our demonstration of aqueous-phase PECD marks a significant advance in the field of liquid-jet photoelectron spectroscopy. Given its chemical-state and site specificity and enantiomeric sensitivity, liquid-phase PECD is likely to become a remarkably useful analytical method for studying chiral molecules under biologically relevant aqueous conditions.
5 Experimental
@@
We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association (HGF), for the provision of experimental facilities. Parts of this research were carried out at PETRA III at the P04 soft x-ray beamline. Beamtime was allocated for proposals II-20180012, I-20200682, II-20210015, I-20211126, and I-20230378. We thank the P04 beamline staff, in particular Dr. Moritz Hoesch and Jörn Seltmann, for their long-standing support for this project. We also thank the European Molecular Biology Laboratory for access to the circular dichroism spectrophotometer. We thank Prof. Petr Slavíček, Prof. Philipp Demekhin, former group members Dr. Anne Stephansen, Dr. Claudia Kolbeck, Dr. Marvin Pohl, Dr. Karen Mudryk, and Dr. Tillmann Buttersack, as well as current group members Harmanjot Kaur, Qi Zhou, Dr. Nicolas Velasquez, Henrik Haak, and Sebastian Kray for valuable discussion and assistance with the transportation and preparation of our setup (EASI) for measurements at PETRA III. We thank Dr. Lukáš Tomaník for support generally and for providing assistance with the representative C 1s measurements specifically.
@@
The supplementary information contains electronic circular dichroism measurements of aqueous solutions of alanine, calculations of the photoionization parameter $b^{0}_{2}$ in the kinetic-energy range of 0–25 eV, a summary of all values of $b^{+1}_{1}$ measured for alanine’s different carbon groups, including values calculated before background subtraction and kinetic-energy averaging was performed, and a description of the means determining the error of the $b^{+1}_{1}$ values, as shown in Figs. 3 and 4.